payton:

What is the solution to the system of equations? y = –3x – 2 5x + 2y = 15 (–40, 19) (–19, 55) (19, –40) (55, –19)

lowkey:

Have you tried solving?

lowkey:

This one is simple as well. You plug in the first equation into the second equation. Like this 5x + 2y = 15 5x + 2( –3x – 2 )= 15

lowkey:

Then you can solve to get x

payton:

distribute?

lowkey:

5x + 2( –3x – 2 )= 15 5x-6x-4=15 -x-4=15 -x=19 x=-19

lowkey:

Yes! and the only answer as -19 as the x value is B

payton:

so b is the obvious answer?

lowkey:

Yes

lowkey:

One way this makes it easier is by going on www.desmos.com then putting in both equations and finding out where the intersection point is (that's the solution)

payton:

its blocked for me

lowkey:

Oh dang. I can keep helping you though!!

payton:

okay thanks, i appreciate it